# Universidad de Granada, Campus de Fuentenueva, E-18071 Granada, Spain

### v1

@@ Abstract
We study constant classical electric fields and the Schwinger effect in de Sitter space, with potential implications for magnetogenesis and inflationary dark matter production. Treating the photon as a dynamical field, we show that sustaining a constant electric field in de Sitter requires a tachyonic photon mass of order the Hubble scale. This observation has physical implications, as it alters the infrared behaviour of the induced Schwinger current. Using an on-shell renormalization condition consistent with a tachyonic photon, we recompute the current for charged fermions and scalars, finding it to be finite and positive even in the massless limit of the charge carriers—contrary to earlier results predicting a puzzling negative IR divergence. For scalars, we include a non-minimal coupling to the Ricci curvature, enabling us to analyze the conformal limit, where the current closely matches that of charged fermions.
@@ 1 Introduction
While such extreme fields remain out of reach in the laboratory, the situation may have been different in the early Universe. During inflation, the vast energy available could naturally generate electric fields strong enough to trigger the Schwinger effect. Moreover, as we discuss in detail, in an expanding spacetime the effect can occur even in the presence of very weak electric fields. As explored in numerous studies Anber:2006xt ; Anber:2009ua ; Graham:2015rva ; Adshead:2016iae ; Bastero-Gil:2018uel ; Bastero-Gil:2021wsf ; Arvanitaki:2021qlj , mechanisms for sustaining a nearly constant electric field during inflation can be constructed by coupling a $U(1)$ gauge field to the background expansion. In this context, the Schwinger effect may play a significant role in both magnetogenesis and the production of dark matter during inflation and raises the intriguing possibility that the first observational evidence of the Schwinger effect could arise from cosmological signals Chua:2018dqh .
@@
The Schwinger current in de Sitter space was first computed in $1+1$ dimensions Frob:2014zka , and later extended to $3+1$ dimensions for minimally coupled charged particles of spin 0 Kobayashi:2014zza ; Hayashinaka:2016dnt ; Banyeres:2018aax and spin 1/2 Hayashinaka:2016qqn . Although the current is formally divergent, it can nevertheless be obtained analytically in a fully non-perturbative manner and made finite through an appropriate renormalization procedure. In $3+1$ dimensions the current was found to display some puzzling features and in particular both the scalar Kobayashi:2014zza ; Banyeres:2018aax and fermion Hayashinaka:2016dnt currents were found to have a peculiar negative infrared (IR) divergence in the small electron  [^2] For the remainder of this paper we use the label “electron” for both charged scalars and fermions. mass regime. This would lead to currents flowing opposite to the electric field direction and imply a negative conductivity. It was argued in Banyeres:2018aax that these regimes with IR negative conductivity are not physical and should be interpreted by taking into account the running of the gauge coupling from the small electron mass scale to the Hubble scale. When the separation between the two scales is very large, one crosses a Landau pole, beyond which the conductivity appears to be negative but the theory is no longer under control.
@@
In this work we argue that the appearance of these IR divergences is due to the renormalization procedures implicitly utilized in Kobayashi:2014zza ; Banyeres:2018aax ; Hayashinaka:2016dnt . Imposing physical renormalization conditions at the Hubble scale we show that the renormalized current is positive in all regimes, both for scalars and fermions, without any need to run the coupling. In deciding how to fix the renormalization conditions, we first point out that in order to sustain a constant electric field in de Sitter space it is unavoidable to have a tachyonic instability. Requiring the electric field to be constant and uniform both inside and outside the Hubble horizon, as was assumed in Kobayashi:2014zza ; Banyeres:2018aax ; Hayashinaka:2016dnt , implies that the photon, treated as a classical dynamical field, must have a tachyonic mass $m_{A}^{2}=-2H^{2}$, where $H$ is the Hubble constant.  Upon quantization of the electron field, the theory contains a single counterterm to absorb the only divergence corresponding to the wave-function renormalization of the photon field or, equivalently, to the electric charge renormalization. We fix the counterterm by imposing an on-shell renormalization condition with $m_{A}^{2}=-2H^{2}$ and show that in doing so one obtains a current which is always positive and free of IR divergences. We also show for the first time that for a conformally coupled scalar the behavior of the current is the same as for a massless fermion, as one would expect.
@@ 2.1 Conditions for a constant electric field
A realistic model could be constructed, for example, by considering a massless photon with an effective coupling to the inflaton $\phi$ such as an axion like coupling $\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$. This acts as a source for the electric field, leads Anber:2009ua ; Bastero-Gil:2018uel to tachyonic production of long wavelength modes for one of the two transverse components of the gauge field, and generates a classical electromagnetic field with $\langle\vec{E}\cdot\vec{B}\rangle\neq 0$, coherent over scales comparable to the Hubble horizon. A full treatment of the Schwinger effect incorporating this mechanism for generating a constant electromagnetic field, including an explicit inflationary model and possible backreaction effects, requires numerical study Sobol:2020frh which we leave for ongoing work followup2 .
@@ 2.2 Renormalized Lagrangian with a constant electric field
In the fully quantum theory of QED there are four quantities (ignoring vacuum energy) which need to be renormalized, associated with the normalization of the fields for the photon and electron, the electron mass, and the electric charge. These four bare parameters receive quantum corrections at loop order which are divergent. However, since we are treating the electric field as classical, in other words the photon is not quantized, there are no internal photon propogators in the theory at any loop order. Thus, the electron propagator is not corrected at loop level which implies that neither the electron field or mass are renormalized. Furthermore, due to the Ward identity, the renormalization of the electric charge is determined by the renormalization of the photon field. So in the end, when considering a classical constant electric field, there is only one Lagrangian parameter which must be renormalized which we take to be the normalization of the photon field.
@@
Varying the action of Eq. (8) with respect to the gauge field we obtain,
@@
2.3 Counterterm and on-shell renormalization conditions
Turning now to the calculation of the counterterm, in principle we should compute the loop correction from electrons to the photon two-point function in de Sitter space. To our knowledge, there are no analytic results currently available for such a calculation. Thus, to be able to compute the counterterm analytically we start by considering the two point function for a massive photon in Landau gauge in flat spacetime,
$\left\langle{A_{\mu}A_{\nu}}\right\rangle=\frac{-i}{p^{2}+m_{A}^{2}}\left(g_{\mu\nu}-\frac{p_{\mu}p_{\nu}}{p^{2}}\right)\,.$ (12)
We need to compute the sum of all one-particle-irreducible (1PI) contributions to Eq. (12), which can be expressed as $\Pi_{\mu\nu}(p)=(p^{2}g_{\mu\nu}-p_{\mu}p_{\nu})\Pi(p^{2})$. As discussed above, solely by computing the one-loop diagram in Fig. 1 we obtain $\Pi(p^{2})$ to all orders in this theory. Since the gauge field is classical this suggests the obvious choice of picking on-shell renormalization conditions. We thus require the two-point function to have a pole at $p^{2}=-m_{A}^{2}$ with residue i-i which then fixes the loop function,
$\Pi(p^{2}=-m_{A}^{2}=2H^{2})=0\,.$ (13)
Writing $\Pi(p^{2})=e^{2}\Pi_{2}(p^{2})+\delta_{3}$ where $\Pi_{2}(p^{2})$ is extracted directly from the diagram of Fig. 1, we obtain the counterterm (to all orders),
$\delta_{3}=-e^{2}\Pi_{2}(2H^{2}).$ (14)
The task now is to compute $\Pi_{2}(p^{2})$ proceeding with the calculation in flat spacetime, but including as much information from the curved background as possible. Consider the equations of motion in a curved background for a free non-minimally coupled scalar,
@@ 2.3 Counterterm and on-shell renormalization conditions
With these ingredients, the calculation of $\Pi_{2}(p^{2})$ in Minkowski space reduces to a classic exercise in quantum field theory. Utilizing Pauli-Villars regularization to control divergences, we impose the renormalization condition in Eq. (14) to obtain the counterterm (see Appendix B.1 for details),
$\delta_{3}^{\phi}$ $=$ $\frac{e^{2}}{144\pi^{2}}\Big{[}-3\ln\frac{\Lambda^{2}}{H^{2}}+3\ln\left(\overline{m}_{\phi}^{2}+12\xi\right)-12\left(\overline{m}_{\phi}^{2}+12\xi\right)$ (17)
$+$ $6\left(2\left(\overline{m}_{\phi}^{2}+12\xi\right)+1\right)^{3/2}\coth^{-1}\left(\sqrt{2\left(\overline{m}_{\phi}^{2}+12\xi\right)+1}\right)-8\Big{]}\,$,
for the scalar case, while for the fermion case we obtain (see Appendix B.2 for details)
$\delta_{3}^{\psi}$ $=$ $\frac{e^{2}}{36\pi^{2}}\Big{[}-3\ln\frac{\Lambda^{2}}{H^{2}}+3\ln\left(\overline{m}_{\psi}^{2}+3\right)+6(\overline{m}_{\psi}^{2}+3)$ (18)
- $6\left((\overline{m}_{\psi}^{2}+3)-1\right)\sqrt{2(\overline{m}_{\psi}^{2}+3)+1}\,\coth^{-1}\left({\sqrt{2(\overline{m}_{\psi}^{2}+3)+1}}\right)-5\Big{]}\,.$
Here $\Lambda$ is the mass of the Pauli-Villars fields and we have taken $R=12H^{2}$ for the curvature in de Sitter space as well as defined $\overline{m}_{\phi,\psi}\equiv m_{\phi,\psi}/H$. We see explicitly that our renormalization condition produces non-trivial finite pieces in the counterterm. They will lead to renormalized currents which are both UV and IR finite as well as always positive. Before continuing with the calculation of the current a few comments are in order:
@@
  • As we examine explicitly below, if we did not include the curvature correction terms in the propagators, the finite terms in $\delta_{3}$ would be different and would lead to a renormalized current which would become negative in some IR regions with $m_{\phi,\psi}/H\ll 1$. This is a justification a posteriori for keeping those terms in the propagators, but it is clear that the full de Sitter calculation of $\delta_{3}$ would take them into account. As we’ll show below, including them in the Minkowski calculation leads to a renormalized current which is always well behaved in the IR and always positive, as expected.

### v2

@@ Abstract
We study constant classical electric fields and the Schwinger effect in de Sitter space, with potential implications for magnetogenesis and inflationary dark matter production. Treating the photon as a dynamical field, we show that sustaining a constant electric field in de Sitter requires a tachyonic photon mass of order the Hubble scale. This observation has physical implications, as it alters the infrared behaviour of the induced Schwinger current. Using an on-shell renormalization condition consistent with a tachyonic photon, we recompute the current for charged fermions and scalars, finding it to be finite and positive even in the massless limit of the charge carriers, contrary to earlier results in the literature predicting a puzzling negative IR divergence. For scalars, we include a non-minimal coupling to the Ricci curvature, enabling us to analyze the conformal limit, where the current closely matches that of charged fermions.
@@ 1 Introduction
While such extreme fields remain out of reach in the laboratory, the situation may have been different in the early Universe. During inflation, the vast energy available could naturally generate electric fields strong enough to trigger the Schwinger effect. Moreover, as we discuss in detail, in an expanding spacetime the Schwinger effect can occur even in the presence of very weak electric fields. As explored in numerous studies Anber:2006xt ; Anber:2009ua ; Graham:2015rva ; Adshead:2016iae ; Bastero-Gil:2018uel ; Bastero-Gil:2021wsf ; Arvanitaki:2021qlj , mechanisms for sustaining a nearly constant electric field during inflation can be constructed by coupling a $U(1)$ gauge field to the background expansion. In this context, the Schwinger effect may play a significant role in both magnetogenesis and the production of dark matter during inflation and raises the intriguing possibility that the first observational evidence of the Schwinger effect could arise from cosmological signals Chua:2018dqh .
@@
The Schwinger current in de Sitter space was first computed in $1+1$ dimensions Frob:2014zka ; Stahl:2015gaa ; Stahl:2016geq , and later extended to $3+1$ dimensions for minimally coupled charged particles of spin 0 Kobayashi:2014zza ; Bavarsad:2016cxh ; Hayashinaka:2016dnt ; Banyeres:2018aax ; Bavarsad:2017oyv ; Stahl:2018idd and spin 1/2 Hayashinaka:2016qqn . Although the regularized current is formally divergent, it can nevertheless be obtained analytically in a fully non-perturbative manner and made finite through an appropriate renormalization procedure. In $3+1$ dimensions the renormalized current was found to display some puzzling features and in particular both the scalar Kobayashi:2014zza ; Banyeres:2018aax and fermion Hayashinaka:2016dnt currents were found to have a peculiar negative infrared (IR) divergence in the small electron  [^2] For the remainder of this paper we use the label “electron” for both charged scalars and fermions. mass regime. This would lead to large currents flowing opposite to the electric field direction and imply a negative conductivity for small electron masses. It was argued in Banyeres:2018aax that these regimes with divergent IR negative conductivity are not physical and should be interpreted by taking into account the running of the gauge coupling from the small electron mass scale to the Hubble scale. When the separation between the two scales is very large, one crosses a Landau pole, beyond which the conductivity appears to be negative but the theory is no longer under control.
@@
In this work we argue that the appearance of these IR divergences is due to the renormalization procedures implicitly utilized in Kobayashi:2014zza ; Banyeres:2018aax ; Hayashinaka:2016dnt . Imposing physical renormalization conditions which are consistent with a constant electric field in de Sitter space, we show that the renormalized current is free of any IR divergences, both for scalars and fermions, without any need to run the coupling. In deciding how to fix the renormalization conditions, we first point out that in order to sustain a constant electric field in de Sitter space it is unavoidable to have a tachyonic instability. Requiring the electric field to be constant and uniform both inside and outside the Hubble horizon, as was assumed in Kobayashi:2014zza ; Banyeres:2018aax ; Hayashinaka:2016dnt , implies that the photon, treated as a classical dynamical field, must have a tachyonic mass $m_{A}^{2}=-2H^{2}$, where $H$ is the Hubble constant. Upon quantization of the electron field, the theory contains a single counterterm to absorb the only divergence corresponding to the wave-function renormalization of the photon field or, equivalently, to the electric charge renormalization. We fix the counterterm by imposing an on-shell renormalization condition with $m_{A}^{2}=-2H^{2}$ and show that one obtains a renormalized current free of UV and IR divergences. We also show for the first time that for a conformally coupled charged scalar the behavior of the current is the same as for a charged fermion, as one would expect.
@@ 2.1 Conditions for a constant electric field
A realistic model could be constructed, for example, by considering a massless photon with an effective coupling to the inflaton $\phi$ such as an axion like coupling $\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$. This acts as a source for the electric field, leads Anber:2009ua ; Bastero-Gil:2018uel to tachyonic production of long wavelength modes for one of the two transverse components of the gauge field, and generates a classical electromagnetic field with $\langle\vec{E}\cdot\vec{B}\rangle\neq 0$, coherent over scales comparable to the Hubble horizon. A full treatment of the Schwinger effect incorporating this mechanism for generating a constant electromagnetic field, including an explicit inflationary model and possible backreaction effects, requires numerical study Sobol:2020frh and left to ongoing work followup2 .
@@ 2.2 Renormalized Lagrangian with a constant electric field
In the fully quantum theory of QED there are four quantities (ignoring vacuum energy) which need to be renormalized, associated with the normalization of the fields for the photon and electron, the electron mass, and the electric charge. These four bare parameters receive quantum corrections at loop order which are divergent. However, since we are treating the electric field as classical, in other words the photon is not quantized, there are no internal photon propagators in the theory at any loop order. Thus, the electron propagator is not corrected at loop level which implies that neither the electron field or mass are renormalized. Furthermore, due to the Ward identity, the renormalization of the electric charge is determined by the renormalization of the photon field. So in the end, when considering a classical constant electric field, there is only one Lagrangian parameter which must be renormalized which we take to be the normalization of the photon field.
@@
In order to define our theory we must first fix the counterterm with an appropriate physical renormalization condition. Since the field $A_{\mu}$ is classical, the obvious choice is to pick on-shell renormalization conditions, requiring the renormalized two point function to have a pole at $p^{2}=-m_{A}^{2}$ with residue i-i which fixes,
$\Pi(p^{2}=-m_{A}^{2})=0,$ (9)
where $e^{2}\Pi(p^{2})$ is defined as the coefficient of $i\left(p^{2}g^{\mu\nu}-p^{\mu}p^{\nu}\right)$ in the sum of all 1PI contributions to the photon 2-point correlation function Schwartz:2014sze . Since we have a classical electric field and $A_{\mu}$ is not quantized, $\Pi(p^{2})$ is determined exactly at one loop, with the corresponding diagram shown in Fig. 1 (with a second diagram for scalars). This renormalization condition leads to the condition on the counterterm,
$\delta_{3}=-e^{2}\Pi(-m_{A}^{2}).$ (10)
Obviously, for a massless photon one takes $m_{A}^{2}=0$ which holds in flat space  [^3] As we discuss more below, this is equivalent to the renormalization procedures used in Kobayashi:2014zza ; Hayashinaka:2016qqn ; Banyeres:2018aax .. However, as we have discussed above, a massless photon is inconsistent with a constant classical electric field in de Sitter space which requires a tachyonic mass $m_{A}^{2}=-2H^{2}$. Thus we take the counterterm to be fixed by the condition,
$\delta_{3}=-e^{2}\Pi(2H^{2}).$ (11)
As we’ll see, with this renormalization condition, the renormalized Schwinger current in de Sitter space is not only UV finite, but crucially, it is also free of the negative IR divergences which were found in previous calculations for both scalars Kobayashi:2014zza ; Banyeres:2018aax and fermions Hayashinaka:2016qqn . In Appendix A we discuss an alternative way of producing a uniform and constant electric field, but no magnetic field, relying on a modified kinetic term. We show that at the end of the day the equation of motion for the vector field is the same as in the tachyonic mass case. That implies that the two cases are physically indistinguishable and that the same renormalization condition described in this section applies to both.
2.3 Counterterm with a constant electric field
Unlike for the regularized Schwinger current, when computing the vacuum polarization diagram we cannot compute analytically the loop diagram in de Sitter space. To obtain an analytic result, one must work in the Minkowski limit in both the spacetime integral and the massive charged scalar propagator, as was done (implicitly) in previous calculations in the literature Banyeres:2018aax . This properly takes care of the UV modes in the counterterm thus allowing divergences in the regularized current to be canceled ensuring a finite renormalized current. The vacuum polarization diagram is then given by the well known result from QED which depends on the invariant mass squared $(p^{2})$ of the photon. If we follow the renormalization procedure used (implicitly) in previous calculations in the literature, assuming a massless photon and setting $p^{2}=0$, a negative IR divergence is introduced into the counterterm when the electron mass is taken to zero. As we’ll see, this negative IR divergence is then introduced into the physical renormalized current as there is no corresponding IR divergence in the regularized current which could cancel it.
To see this explicitly we take the flat space result from scalar QED for the vacuum polarization Schwartz:2014sze . Setting $p^{2}=0$, we have for the counterterm,
$\delta_{3}^{\phi}$ $=$ $-e^{2}\Pi(0)=\left(\frac{e}{12\pi}\right)^{2}\Big[-3\ln\Big(\frac{\Lambda^{2}}{m_{\phi}^{2}}\Big)\Big]\,$, (12)
where $m_{\phi}$ is the mass of the charged scalar which enters through the propogator (in the Minkowski limit) and $\Lambda$ is the mass of the Pauli-Villars fields used to regulate the UV divergence (see Appendix B.1 for details) which cancels the one in the regularized current. With this counterterm we can reproduce previous results in the literature for the renormalized current of a minimally coupled massive charged scalar Kobayashi:2014zza ; Banyeres:2018aax which were obtained using different renormalization procedures. While the UV log divergence cancels the one in the regularized current we see explicitly the IR divergence in the $m_{\phi}\to 0$ limit. Since the vacuum polarization is computed strictly in the Minkowski limit, it knows nothing about de Sitter and is independent of $H$ which we expect plays a role in the infrared part of the calculation, as we’ll see explicitly below. Note also that one cannot invoke renormalization of the electric charge to simply throw away the $\ln(m_{\phi}^{2}/H^{2})$ term, as advocated in Banyeres:2018aax , since this term is needed to ensure the proper behavior of the renormalized current in the large $m_{\phi}/H$ limit, as we’ll see explicitly in Section 3.
However, as we have discussed above, taking $p^{2}=0$ is inconsistent with the presence of a constant (dynamical) electric field in de Sitter space, independently of considerations about renormalization or the Schwinger current. As we have shown, in order to sustain a constant (dynamical) electric field in de Sitter space one requires the photon to have a tachyonic mass $p^{2}=m_{A}^{2}=-2H^{2}$. If we evaluate the vacuum polarization diagram now with $p^{2}=2H^{2}$ and again in the Minkowski limit we instead obtain,
$\delta_{3}^{\phi}$ $=$ $-e^{2}\Pi(2H^{2})=\left(\frac{e}{12\pi}\right)^{2}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)+3\ln\Big(\frac{m_{\phi}^{2}}{H^{2}}\Big)$ (13)
$+$ $6\Big(2\frac{m_{\phi}^{2}}{H^{2}}+1\Big)^{3/2}\coth^{-1}\Big(\sqrt{2\frac{m_{\phi}^{2}}{H^{2}}+1}\Big)-12\frac{m_{\phi}^{2}}{H^{2}}-8\Big]\,$,~~~~~~~~
where we see explicitly in the second line the new terms which are now generated. Crucially, in the limit $m_{\phi}/H\to 0$ there is no negative IR divergence and $\delta_{3}^{\phi}$ converges to,
$\delta_{3}^{\phi}$ $=$ $\frac{e^{2}}{144\pi^{2}}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)+\ln 8-8\Big]\approx\frac{e^{2}}{144\pi^{2}}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)-5.9\Big].$ (14)
The log divergence cancels the one in the regularized current giving a renormalized current which is both UV and IR finite in the massless electron limit.
Repeating the analysis for a charged fermion loop we again first examine the counterterm assuming a massless photon with $p^{2}=0$,
$\delta_{3}^{\psi}$ $=$ $-e^{2}\Pi(0)=\left(\frac{e}{6\pi}\right)^{2}\Big[-3\ln\Big(\frac{\Lambda^{2}}{m_{\psi}^{2}}\Big)\Big]\,$, (15)
where $m_{\psi}$ is the charged fermion mass and again we see the IR divergence in the $m_{\psi}\to 0$ limit. With this counterterm one reproduces the results for the renormalized current from Hayashinaka:2016qqn which computed the current using adiabatic subtraction. Now imposing the renormalization condition in Eq. (11) consistent with a constant electric field in de Sitter space we obtain,
$\delta_{3}^{\psi}$ $=$ $-e^{2}\Pi(2H^{2})=\frac{e^{2}}{36\pi^{2}}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)+3\ln\Big(\frac{m_{\psi}^{2}}{H^{2}}\Big)$
- $6\Big(\frac{m_{\psi}^{2}}{H^{2}}-1\Big)\sqrt{2\frac{m_{\psi}^{2}}{H^{2}}+1}\,\coth^{-1}\Big({\sqrt{2\frac{m_{\psi}^{2}}{H^{2}}+1}}\Big)+6\frac{m_{\psi}^{2}}{H^{2}}-5\Big],$
where again in the second line we see the new terms generated by the tachyonic photon mass condition. Again this is free of IR divergences in the $m_{\psi}/H\to 0$ limit which gives,
$\delta_{3}^{\psi}$ $=$ $\frac{e^{2}}{36\pi^{2}}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)+\ln 8-5\Big]\approx\frac{e^{2}}{36\pi^{2}}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)-2.9\Big].$ (17)
We also see that the log divergent and log 8 terms have the same coefficient as the scalar case up to a factor of 4. As in the scalar case, the log divergence cancels the one in the regularized current giving a renormalized fermion current which is both UV and IR finite in the massless electron limit. While Eq. (14) and Eq. (17) are obtained in the Minkowski and large loop momentum limit, they are perfectly self consistent and free of IR divergences which have been cured, not by hand, but by imposing the tachyonic photon mass condition needed for sustaining a constant electric field in de Sitter space.
2.4 Renormalized Schwinger current and modified counterterms
Until now, our previous discussion has been completely independent of considerations of the Schwinger current. With the counterterm defined in Eq. (11) we can define the physical renormalized current. Varying the action of Eq. (8) with respect to the gauge field,
@@ 2.4 Renormalized Schwinger current and modified counterterms
The regularized current is obtained non-perturbatively containing all of the effects coming from de Sitter space and full dependence on the electron mass where crucially, there is no IR divergence in the small electron mass limit. However, the counterterm in Eq. (14) has been obtained taking the Minkowski limit in the vacuum polarization diagram. Thus, although there is no IR divergence when taking $p^{2}=2H^{2}$, we should not expect to properly capture all IR effects. We can see this when considering the renormalized Schwinger current where, because of the missing contributions from de Sitter space, this procedure for computing the vacuum polarization diagram leads to a small (constant) negative finite contribution to the counterterm of roughly $\delta_{3\rm{fin}}^{\phi}\approx-0.004e^{2}$ in Eq. (14) for a charged scalar and $\delta_{3\rm{fin}}^{\psi}\approx-0.008e^{2}$ in Eq. (17) for a charged fermion. As we’ll see below, in the small electron mass limit, this leads to a constant, but slightly negative renormalized current implying the unphysical behavior of a current flowing opposite to the electric field. As we explore more below, we expect that including additional corrections to the vacuum polarization from de Sitter space ensures a positive current.
In general corrections from de Sitter space cannot be included analytically, but one that can be included analytically is the curvature correction to the electron mass coming from a non-minimal coupling to gravity in the case of a scalar or the spin connection as in the case of the charged fermion. These corrections are automatically included in the regularized current which is computed non-perturbatively in de Sitter space directly from the Lagrangian. However, in the counterterm they are not included in the pure Minkowski approximation of the electron propagator used above and implicitly in previous calculations in the literature. To see how they enter in the electron propagator we can consider the equations of motion in a curved background for a free non-minimally coupled scalar,
@@
With the modified propagators we repeat the calculation of the vacuum polarization imposing the tachyonic photon mass in the renormalization condition in Eq. (11) to obtain the counterterm (see Appendix B.1 for details),
$\delta_{3}^{\phi}$ $=$ $\frac{e^{2}}{144\pi^{2}}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)+3\ln\Big(\frac{m_{\phi}^{2}}{H^{2}}+12\xi\Big)-12\Big(\frac{m_{\phi}^{2}}{H^{2}}+12\xi\Big)$ (23)
$+$ $6\Big(2\Big(\frac{m_{\phi}^{2}}{H^{2}}+12\xi\Big)+1\Big)^{3/2}\coth^{-1}\Big(\sqrt{2\Big(\frac{m_{\phi}^{2}}{H^{2}}+12\xi\Big)+1}\Big)-8\Big]\,$,
for the scalar case, while for the fermion case we obtain (see Appendix B.2 for details),
$\delta_{3}^{\psi}$ $=$ $\frac{e^{2}}{36\pi^{2}}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)+3\ln\Big(\frac{m_{\psi}^{2}}{H^{2}}+3\Big)+6\Big(\frac{m_{\psi}^{2}}{H^{2}}+3\Big)$ (24)
- $6\Big(\Big(\frac{m_{\psi}^{2}}{H^{2}}+3\Big)-1\Big)\sqrt{2\Big(\frac{m_{\psi}^{2}}{H^{2}}+3\Big)+1}\,\coth^{-1}\Big({\sqrt{2\Big(\frac{m_{\psi}^{2}}{H^{2}}+3\Big)+1}}\Big)-5\Big]\,$,
where again $\Lambda$ is the mass of the Pauli-Villars fields and we have taken $R=12H^{2}$ for the curvature in de Sitter space. We see explicitly that the curvature corrections to the mass give additional positive finite contributions in the counterterm. Now in the massless limit,
$\delta_{3}^{\phi}$ $=$ $\frac{e^{2}}{144\pi^{2}}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)+\ln 8-32+30\sqrt{5}\cot^{-1}(\sqrt{5})\Big]$ (25)
$\approx$ $\frac{e^{2}}{144\pi^{2}}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)+2.4\Big]\,$,
where we have taken $\xi=1/6$ for the case of a conformally coupled scalar. So we see explicitly the positive contributions coming from the curvature corrections to the charged scalar mass ensure a positive finite contribution [^4] Note for the minimally coupled scalar case with $\xi=0$ there is no curvature correction to the mass. In this case the finite contribution from the counterterm is negative as in Eq. (14), but the regularized current is positive definite with a larger magnitude than the negative finite piece of the counterterm so the renormalized current is always positive in this case as well (see solid curves in Fig. 2). to the counterterm of $\delta_{3\rm{fin}}^{\phi}\approx 0.002e^{2}$. As we’ll see below, this ensures a positive renormalized scalar current. For the charged fermion counterterm in the massless limit we obtain (see Appendix B.2 for details),
$\delta_{3}^{\psi}$ $=$ $\frac{e^{2}}{36\pi^{2}}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)+\ln(27)+13-12\sqrt{7}\coth^{-1}(\sqrt{7})\Big]$
$\approx$ $\frac{e^{2}}{36\pi^{2}}\Big[-3\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)+3.7\Big],$
which leads to a positive finite contribution to the counterterm of $\delta_{3\rm{fin}}^{\psi}\approx 0.01e^{2}$ and again to a positive renormalized fermion current, as we’ll see below. These counterterms will lead to renormalized currents which are both UV and IR finite as well as always positive.
Before continuing with the calculation we summarize our previous discussion:
@@
  • As we examine explicitly below, if we did not include the curvature correction terms in the propagators, the finite term in $\delta_{3}$ is negative and leads to a renormalized current which would become negative in some IR regions with $m_{\phi,\psi}/H\ll 1$. This is a justification a posteriori for keeping those terms in the propagators, but it is clear that the full de Sitter calculation of $\delta_{3}$ would take them into account. As we’ll show below, including them in the Minkowski calculation leads to a renormalized current which is always well behaved in the IR and always positive, as expected.